OpenStudy (anonymous):

How to factor (-2e^(-2sa)+e^(-sa)+1?

OpenStudy (anonymous):

-2e^(-2sa)+e^(-sa)+1 = -2e^(-2sa)+2e^(-sa)-e^(-sa)+1 = -2e^(-sa)(e^-(sa)+1)-1(e^-(sa)+1) = (e^-(sa)+1)(-2e^-(sa)-1)